A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
Answer by stanbon(48558)

A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
$9 coffee DATA:
amount= x lbs ; value=9x dollars
------------------------------
$12 coffee DATA:
amount: 100-x lbs ; value = 12(100-x)=1200-12x dollars
------------------------------------
Mixed coffee DATA:
amount: 100 lbs ; value- 11.25(100)= 1125 dollars
-------------------------------------
EQUATION:
value + value = $1125
9x + 1200-12x = 1125
-3x=-75
x=25 lbs (amount of $9 coffee)
100-x=75 lbs (amount of $12 coffee)
-----------------------------------
Cheers,
Stan H.